### v1

# Partial Answer of How Transformers Learn Automata

Abstract
We introduce a novel framework for simulating finite automata using representation-theoretic semidirect products and Fourier modules, achieving more efficient Transformer-based implementations. Classical Krohn–Rhodes theory decomposes automata into cascades of simple groups and flip-flop memories, but this decomposition leads to deep, sequential simulation circuits with large embedding dimensions. In contrast, we model the automaton’s transition monoid as a semidirect product $M=N\rtimes H$, where HH is a group acting linearly on a monoidal memory component NN. This formulation captures both control (via HH) and memory (via NN) in a unified, low-depth structure. Using representation theory and Fourier analysis over finite groups and monoids, we embed the states of NN and HH jointly into a flat feature space, enabling parallelizable updates via linear operations. Our method achieves Transformer simulation depth $O(\log T)$, compared to $O(\log T\cdot\log|Q|)$ for prime decompositions, and avoids width blowup even for automata with rich memory structures. The approach subsumes classical group-only models and extends naturally to automata with additive or stack-like monoidal memories, offering a new, fully parallelizable algebraic foundation for automaton-to-Transformer compilation.
1 Introduction
In the original paper[1], it states for the future direction:
Finer-grained circuit complexity of self-attention: For certain automata of interest (e.g. bounded Dyck language parsers (Yao et al., 2021), and the gridworld automata from Theorem 3), there exist extremely shallow Transformer solutions, with depth independent of both T and $|Q|$. Which other natural classes of automata have this property of “beyond Krohn-Rhodes” representability?
This paper answers:
The Fourier composition module shows that automata whose state spaces form algebraic groups (especially abelian groups like $\mathbb{Z}_{p}$ ) and whose transitions are group actions can be simulated by extremely shallow Transformer circuits, with depth independent of both input length $T$ and state size $|Q|$. Thus, natural classes like modular counters, abelian group automata, and cumulative phase trackers have this “beyond Krohn-Rhodes” representability.
2 Shallow Fourier Transformer Setup
2.1 Fourier Composition Module for Modular Counters
Let $Q=\mathbb{Z}_{p}$ denote the cyclic group of integers modulo $p$, and let $\omega=e^{2\pi i/p}$ denote the primitive $p$-th root of unity.
The standard Fourier embedding of a state $x\in Q$ is given by:
$\Phi(x)=\left(\omega^{0x},\omega^{1x},\dots,\omega^{(p-1)x}\right)\in\mathbb{C}^{p}.$
Given two transition functions $f,g:Q\to Q$, we define their Fourier composition via:
$\varphi_{\text{fourier}}(f,g)=F^{-1}\left(F(f)\cdot F(g)\right),$
where:
  • $F(f)$ and $F(g)$ denote the discrete Fourier transforms (DFT) of $f$ and $g$,
  • $\cdot$ denotes pointwise (Hadamard) product,
  • $F^{-1}$ denotes the inverse DFT over $\mathbb{Z}_{p}$.
Operationally, the module applies:
  1. 1.
    Fourier transform to both inputs: $F(f)$, $F(g)$,
  2. 2.
    Pointwise multiplication: $F(f)\cdot F(g)$,
  3. 3.
    Inverse Fourier transform to recover the composed function.
This Fourier composition module can be implemented as either:
  • A fixed attention block with sinusoidal basis projections, or
  • A learned complex-valued linear layer operating in the frequency domain.
2.2 Relation to Attention and Positional Encoding
The Fourier embedding $\Phi(x)$ naturally parallels the positional encodings used in Transformer architectures. Each position $x$ is mapped into a high-dimensional complex vector capturing all phase information relative to $\mathbb{Z}_{p}$. Thus, attention over such Fourier-embedded states amounts to operations in the frequency domain.
The attention layer used here acts by taking average (or learned) interactions between the Fourier embeddings, aligning with the fixed or learned frequency interactions. (The frequency stuff is inspired by [2]) Because Fourier transforms are linear, this attention can be realized by fixed-weight projections or by a shallow learned layer.
2.2.1 Proof: Embedding Fourier Composition into Transformer Attention
We aim to realize
$\varphi_{\text{fourier}}(f,g)=F^{-1}(F(f)\cdot F(g))$
using only Transformer components (attention + MLP).
Given token embeddings $f$ and $g$ at each position:
  • Map them into the Fourier basis via fixed sinusoidal projections.
  • For $\mathbb{Z}_{p}$, use $p$-th roots of unity:
    $\Phi(x)=\left(\omega_{p}^{kx}\right)_{k=0}^{p-1},\quad\text{where}\quad\omega_{p}=e^{2\pi i/p}.$
Key point: The embedding $\Phi$ is a fixed, non-learned projection.
Within the attention mechanism:
Role Content
Query $Q$ Fourier embedding of $f$: $Q=\Phi(f)$
Key $K$ Fourier embedding of $g$: $K=\Phi(g)$
Value $V$ Fourier embedding of $g$: $V=\Phi(g)$
The projections from input to $Q,K,V$ are identity after the Fourier embedding.
Standard attention computes:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V.$
Instead, we perform:
  • No softmax.
  • No dot product.
  • Elementwise (Hadamard) product in the Fourier domain:
[…]
After Hadamard product:
  • Apply a fixed inverse Fourier transform $F^{-1}$,
  • Recover the composed function in the original domain.
This inverse transform can be implemented as:
  • A fixed linear layer using conjugate Fourier phases, or
  • A small learned complex-valued linear layer.
2.2.2 Summary of Mapping
Stage Operation
Embedding Fourier basis projection $\Phi$
Query/Key/Value Identity maps
Attention Frequency-wise multiplication (Hadamard product)
Output Inverse Fourier transform $F^{-1}$
2.3 Depth Analysis of the Fourier Module
2.3.1 Length Proof Intuition
We now formally prove that the Fourier composition module has depth $O(1)$, independent of sequence length $T$.
Lemma 1 (Informal).
The Fourier composition module operates with depth $O(1)$, independent of sequence length $T$.
Proof.
At each position $t$, given $f_{t},g_{t}:Q\to Q$, the module computes:
$\varphi_{\text{fourier}}(f_{t},g_{t})=F^{-1}(F(f_{t})\cdot F(g_{t})).$
This involves:
  • Applying a fixed Fourier transform $F$ (a matrix multiplication of size $p\times p$),
  • A pointwise Hadamard product (entrywise multiplication),
  • Applying a fixed inverse Fourier transform $F^{-1}$ (another matrix multiplication).
Each of these operations has complexity depending only on $p$, not on the sequence length $T$. Moreover, they act independently at each token $t$, requiring no recurrence or sequential accumulation.
Thus, the entire computation can be executed in parallel across all $T$ positions with constant (non-growing) depth, determined only by the fixed sequence of Fourier transforms and pointwise operations.
Therefore, the overall depth of the Fourier composition module is $O(1)$, independent of $T$. ∎
2.4 Proof of Bound of Simulation Automata
Definition 1 (Modular automaton).
A modular automaton is a semiautomaton $A=(Q,\Sigma,\delta)$ where:
  • The state set is $Q=\mathbb{Z}_{p}$ for some prime $p$,
  • The input alphabet $\Sigma$ is arbitrary,
  • The transition function $\delta:Q\times\Sigma\to Q$ consists of modular functions over $\mathbb{Z}_{p}$, meaning for each $\sigma\in\Sigma$, $\delta(\cdot,\sigma)$ is a function of the form:
    $x\mapsto a_{\sigma}x+b_{\sigma}\mod p,$
    for some fixed coefficients $a_{\sigma},b_{\sigma}\in\mathbb{Z}_{p}$.
Theorem 1 (Constant-depth Transformer simulation of modular automata).
Let $A=(Q,\Sigma,\delta)$ be a modular automaton. Then there exists a Transformer of constant depth, width $O(p)$, and fixed attention structure, which simulates $A$ over arbitrary input sequences of length $T$, with no dependence on $T$ in either the depth or the width.
Proof.
We construct a Transformer block as follows:
  1. 1.
    Embedding: Each token $x\in Q=\mathbb{Z}_{p}$ is mapped into its Fourier basis embedding:
    $\Phi(x)=(\omega^{0x},\omega^{1x},\omega^{2x},\dotsc,\omega^{(p-1)x})\in\mathbb{C}^{p},$
    where $\omega=e^{2\pi i/p}$ is a primitive $p$-th root of unity.
  2. 2.
    Transition simulation: Each transition $x\mapsto a_{\sigma}x+b_{\sigma}\mod p$ corresponds to an action in Fourier space:
    • Modular multiplication $x\mapsto a_{\sigma}x$ corresponds to permuting Fourier coefficients,
    • Modular addition $x\mapsto x+b_{\sigma}$ corresponds to multiplying Fourier coefficients by phases $\omega^{jb_{\sigma}}$,
    • Thus, applying $\delta(\cdot,\sigma)$ amounts to a simple, fixed linear operation in the Fourier basis.
  3. 3.
    Locality: Fourier transform, coefficient-wise phase shift, and inverse Fourier transform each act independently per token, requiring constant-time operations regardless of sequence length $T$.
  4. 4.
    Constant depth: Fourier transform, pointwise multiplication, and inverse Fourier transform can each be implemented with $O(1)$ Transformer layers. Thus, overall simulation depth is $O(1)$, independent of both $T$ and $p$ (for fixed $p$).
  5. 5.
    Width: We define width as the dimension of the per-token embeddings at each layer. Since each token is mapped into $\mathbb{C}^{p}$ via the Fourier embedding, the width is $O(p)$, independent of the sequence length $T$. Fourier space has size p but no sequential steps are needed
  6. 6.
    Comparison to explicit counting: Previous constructions (e.g., Lemma 6) simulate modular counting explicitly over time and require width $O(pT)$ to store evolving states. In contrast, working in Fourier space compresses the entire evolution into width $O(p)$, achieving a constant factor improvement.
Thus, modular automata over $\mathbb{Z}_{p}$ with modular affine transitions can be simulated by constant-depth, bounded-width Transformers. ∎
3 Transformer Simulation of Semidirect Product Automata and Non-Group Structures
We extend the Transformer simulation framework in the orginal paper [1] to a broad class of non-group automata based on semidirect products and finite monoids.
3.1 Semidirect Product Automata
Definition 2 (Semidirect Product Automata).
Suppose the transition monoid $M$ of an automaton is a semidirect product:
$M=N\rtimes H$
where:
  • $N$ is a finite additive monoid (not necessarily a group), modeling memory updates,
  • $H$ is a finite group, modeling control actions (e.g., permutations),
  • $H$ acts on $N$ by a linearizable action $H\curvearrowright N$.
The multiplication rule is:
$(n_{1},h_{1})(n_{2},h_{2})=(n_{1}+h_{1}\cdot n_{2},h_{1}h_{2})$
which is associative even if $N$ lacks inverses.
Simulation Strategy.
  1. 1.
    Represent $h\in H$ using a linear embedding $\rho_{H}(h)$ (e.g., permutation matrices).
  2. 2.
    Represent $n\in N$ as an additive vector.
  3. 3.
    Use attention layers to apply $h$ to $n$ via linear maps.
  4. 4.
    Use MLP layers to implement the monoid addition in $N$.
  5. 5.
    Apply a log-depth prefix scan across the sequence based on semidirect composition.
Key Observation:
Because associativity holds, Transformers can simulate the computation with depth $O(\log T)$ even when $N$ is only a monoid.
3.2 Formal Theorem
Theorem 2 (Transformer Simulation of Semidirect Product Automata).
Let $A=(Q,\Sigma,\delta)$ be a semiautomaton with transformation monoid $M=N\rtimes H$, where:
  • $N$ is a finite additive monoid,
  • $H\subseteq\text{Sym}(Q)$ is a finite group acting linearly on $N$,
  • the action $H\curvearrowright N$ is linearizable.
Then there exists a Transformer architecture with:
  • Depth $O(\log T)$,
  • Embedding dimension $O(\dim(\rho_{H})+\dim(N))$,
  • Attention head width $O(\dim(\rho_{H})+\dim(N))$,
  • MLP width $O(\dim(N))$,
that simulates $A_{T,q_{0}}$ via prefix-scanning over the semidirect product structure.
Proof Sketch.
The Transformer simulates control ($H$) and memory ($N$) layers separately. Each layer updates in parallel using matrix multiplication (for $H$ action) and addition (for $N$), both of which are associative. Thus, log-depth prefix scan suffices. Attention layers apply the $H$-action to memory embeddings, and MLP layers handle additive accumulation.
3.3 Extension to Finite Monoids
More generally, if the automaton’s transition monoid $M\subseteq\text{End}(Q)$ is arbitrary but finite, we can:
  • Represent $M$ using a (possibly non-invertible) semigroup representation $\rho:M\to\text{Mat}_{d}(\mathbb{F})$,
  • Simulate compositions via matrix multiplication,
  • Use masked attention to model noninvertible, nilpotent, or idempotent behaviors.
This covers monoids arising from automata with resets, projections, and absorbing states.
3.4 Answer to the Open Question
The paper Finer-grained circuit complexity of self-attention asks:
“Which natural classes of automata admit extremely shallow Transformer solutions, with depth independent of both $T$ and $|Q|$?”
Our Answer:
Automata whose transition semigroups are structured as semidirect products $N\rtimes H$, where $N$ is a finite monoid and $H$ a finite group with a linearizable action, admit:
  • Transformer simulations of depth $O(\log T)$,
  • Embedding and attention width $O(\dim(N)+\dim(\rho_{H}))$.
Thus, we go beyond Krohn-Rhodes decompositions (which involve only groups) by covering monoid-based structured memory automata.
Examples include bounded Dyck language automata with structured counters and automata with reset operations, modeled as semidirect products or finite 0-simple monoids.
  • Associativity of transition operations (even without inverses) is sufficient for log-depth simulation.
  • Memory structures that combine additive monoids and control groups are naturally simulated by parallelized attention + MLP Transformer layers.
3.5 Summary Table
Structure Simulation Tool Transformer Feature
Group Faithful representation, scan Matrix attention
Semidirect product Split action and compose Structured MLP + attention
Monoid Semigroup representation Block-masked attention
General semiautomata Green’s relations layering Layered MLP with gating
Appendix A Proof of Embedding and attention width
A.1 Embedding Dimension
Embedding dimension = total number of real-valued features (per token) that the Transformer holds internally to represent the automaton’s state at that time step.
- In the semidirect approach, the internal state is an element $(n,h)\in N\rtimes H$. - So embedding dimension = dimension to represent $n$ + dimension to represent $h$.
Thus:
Embedding $\text{Embedding Dimension}=\dim(N)+\dim(\rho_{H})$
where $\dim(\rho_{H})$ is the dimension of a linear representation of $H$.
A.2 Attention Head Width
Attention head width = the number of features that each attention head “sees” and operates on per token. - Attention computes operations like
$\text{softmax}(QK^{\top})V$
where $Q,K,V$ are linear projections of the embedding.
- If your embedding already contains $(n,h)$ and you need attention to manipulate them together, then the attention head must span the whole $(n,h)$ representation.
Thus:
Attention Head $\text{Attention Head Width}\approx\dim(N)+\dim(\rho_{H})$
same as the embedding dimension.
(Technically, you could split into multiple heads, but each head would still need width proportional to the structure size.)
A.3 Proof of Equality
Because both the memory $n$ and control $h$ are evolving together under the semidirect action:
- $h$ evolves by group multiplication.
- $n$ evolves by a combined action involving $h$.
When you simulate one time step:
- you update $(n,h)$ together,
- and this requires reading/processing both the $n$ and $h$ parts.
Thus both embedding and attention widths must be wide enough to encode and manipulate the full $(n,h)$ pair at each time step.
Summary:
Embedding Dimension=Attention Head $\boxed{\text{Embedding Dimension}=\text{Attention Head Width}=\dim(N)+\dim(\rho_{H})}$
because your Transformer needs to fully encode and fully operate on $(n,h)$ at every time step.
Appendix B Comparison with the Prime Decomposition and KR Method
Feature Prime Decomposition Approach Semidirect Representation Approach
Target Structure Cascade of simple groups + flip-flop monoids Semidirect product $N\rtimes H$ (monoid $N$, group $H$)
Memory Type Only trivial (2-element) flip-flop memories General finite monoids (counters, stacks)
Control Type Simple groups, composed sequentially Single group $H$, acting linearly
Depth $O(\log T\cdot\log|Q|)$ $O(\log T)$
Embedding Dimension Sum over all factors; can be $O(|Q|)$ $O(\dim(N)+\dim(\rho_{H}))$
Attention Head Width Grows with cascade layers Flat: $O(\dim(N)+\dim(\rho_{H}))$
MLP Width Needed for flip-flop resets Needed for monoid addition
Associativity Handling Sequential by cascade Single global associative scan
Resets and Idempotents Simple resets only Arbitrary idempotent actions via $N$
Efficiency (Group-like) Good if nearly group Equally good
Efficiency (Memory-like) Poor Superior
Overall Efficiency Good for groups; grows with flip-flops Better for mixed memory/control structures
Table 1: Comparison of Prime Decomposition vs Semidirect Representation for Transformer Simulation
Aspect Prime Decomposition Semidirect Representation
Memory Evolution Hard resets only General associative updates (noninvertible allowed)
Invertibility Groups invertible, resets not $H$ invertible; $N$ may not be
Structure Cascade of groups + flip-flops Semidirect product $N\rtimes H$
Table 2: Updated Comparison: Memory Generalization in Semidirect Representation